# On geometric spectral functionals

### v1

@@ 4.2.1 The Euler and Hodge gradings
$\chi_{e}=\gamma^{1}\gamma^{2}\cdots\gamma^{n}\tilde{\gamma}^{1}\tilde{\gamma}^{2}\cdots\tilde{\gamma}^{n},\qquad$ (Euler grading) (4.32)
$\chi_{h}=i^{m}\gamma^{1}\gamma^{2}\cdots\gamma^{n},\qquad$ (Hodge grading)
It could be easily verified that on a form of degree $p$ on a manifold of dimension $2m$, $\chi_{e}$ is $(-1)^{p}$ and $\chi_{h}$ is simply $i^{p(p-1)-m}\ast$, where $\ast$ is the Hodge map. Note that the composition of both gradings is another grading, $\hat{\chi}=\chi_{h}\chi_{e}=i^{m}\tilde{\gamma}^{1}\tilde{\gamma}^{2}\cdots\tilde{\gamma}^{n}$, which, however, commutes with $d+d^{\ast}$. Furthermore, note that
$\chi_{h}\lambda_{\pm}^{p}=\lambda_{\mp}^{p}\chi_{h},\qquad\hat{\chi}\lambda_{\pm}^{p}=-\lambda_{\mp}^{p}\hat{\chi}.$ (4.33)
These gradings behave quite differently and, for example, can restrict perturbations of the Dirac operator. While the most general torsion term anticommutes with $\chi_{e}=\chi_{h}\hat{\chi}$, we obtain a nontrivial condition for the anticommutation of the torsion term with $\chi_{h}$ alone, $T_{ijk}(\lambda^{j}_{+}\lambda^{i}_{+}\lambda^{k}_{-}+\lambda^{k}_{+}\lambda^{i}_{-}\lambda^{j}_{-})=-T_{ijk}(\lambda^{j}_{-}\lambda^{i}_{-}\lambda^{k}_{+}+\lambda^{k}_{-}\lambda^{i}_{+}\lambda^{j}_{+})$, leading to $T_{ijk}(\delta^{ik}\lambda^{j}_{+}-\delta^{jk}\lambda^{i}_{+})=0$, so that $T_{ijj}=0$.
@@
The chiral metric functionals for the Hodge-Dirac operator with torsion, for the gradings $\chi_{h}$, $\chi_{e}$, and $\hat{\chi}$, are given by
$\mathscr{g}_{D_{0}}^{\chi_{h}}(u,w)=\begin{cases}-8\pi i\int_{M}\mathrm{dvol}_{g}u_{a}w_{b}\varepsilon^{ab},\quad n=2,\\ 0,\quad n>2,\end{cases}$ (4.34)
$\mathscr{g}_{D_{0}}^{\chi_{e}}(u,w)=0,\quad n\geq 2,$ (4.35)
@@
It is worth noting that the chiral torsion and Einstein functionals are, in principle, explicitly computable using Proposition 4.4 and Proposition 4.2, respectively. These calculations are not performed in this study, as they do not significantly contribute to the objectives of the current research. Interested readers may pursue the details independently if they wish.
@@ Acknowledgements
The work of AB was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2111 – 390814868. AB is supported by the Alexander von Humboldt Foundation. A.S. is supported by the Polish National Science Centre grant 2020/37/B/ST1/01540. L.D. is affiliated with GNFM–INDAM (Istituto Nazionale di Alta Matematica) and acknowledges that this research is part of the EU Staff Exchange project 101086394 "Operator Algebras That One Can See".

### v2

@@ 4.2.1 The Euler and Hodge gradings
$\chi_{e}=\gamma^{1}\gamma^{2}\cdots\gamma^{n}\tilde{\gamma}^{1}\tilde{\gamma}^{2}\cdots\tilde{\gamma}^{n}\qquad\hbox{(Euler grading)},$
$\chi_{h}={\rm i}^{m}\gamma^{1}\gamma^{2}\cdots\gamma^{n}\qquad\hbox{(Hodge grading)}.$
It could be easily verified that on a form of degree $p$ on a manifold of dimension $2m$, $\chi_{e}$ is $(-1)^{p}$ and $\chi_{h}$ is simply ${\rm i}^{p(p-1)-m}\ast$, where $\ast$ is the Hodge map. Note that the composition of both gradings is another grading, $\hat{\chi}=\chi_{h}\chi_{e}={\rm i}^{m}\tilde{\gamma}^{1}\tilde{\gamma}^{2}\cdots\tilde{\gamma}^{n}$, which, however, commutes with $d+d^{\ast}$. Furthermore, note that $\chi_{h}\lambda_{\pm}^{p}=\lambda_{\mp}^{p}\chi_{h}$, $\hat{\chi}\lambda_{\pm}^{p}=-\lambda_{\mp}^{p}\hat{\chi}$.
These gradings behave quite differently and, for example, can restrict perturbations of the Dirac operator. While the most general torsion term anticommutes with $\chi_{e}=\chi_{h}\hat{\chi}$, we obtain a nontrivial condition for the anticommutation of the torsion term with $\chi_{h}$ alone, $T_{ijk}\bigl(\lambda^{j}_{+}\lambda^{i}_{+}\lambda^{k}_{-}+\lambda^{k}_{+}\lambda^{i}_{-}\lambda^{j}_{-}\bigr)=-T_{ijk}\bigl(\lambda^{j}_{-}\lambda^{i}_{-}\lambda^{k}_{+}+\lambda^{k}_{-}\lambda^{i}_{+}\lambda^{j}_{+}\bigr)$, leading to $T_{ijk}\bigl(\delta^{ik}\lambda^{j}_{+}-\delta^{jk}\lambda^{i}_{+}\bigr)=0$, so that $T_{ijj}=0$.
@@
The chiral metric functionals for the Hodge–Dirac operator with torsion, for the gradings $\chi_{h}$, $\chi_{e}$, and $\hat{\chi}$, are given by
$\mathscr{g}_{D_{0}}^{\chi_{h}}(u,w)=\begin{cases}\displaystyle-8\pi{\rm i}\int_{M}\mathrm{dvol}_{g}u_{a}w_{b}\varepsilon^{ab},&n=2,\\ 0,&n>2,\end{cases}\qquad\mathscr{g}_{D_{0}}^{\chi_{e}}(u,w)=0,\qquad n\geq 2,$
@@
It is worth noting that the chiral torsion and Einstein functionals are, in principle, explicitly computable using Propositions 4.3 and 4.2, respectively. These calculations are not performed in this study, as they do not significantly contribute to the objectives of the current research. Interested readers may pursue the details independently if they wish.
@@ Acknowledgements
We thank the anonymous referees for their insightful comments. The work of AB was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868. AB is supported by the Alexander von Humboldt Foundation. A.S. is supported by the Polish National Science Centre grant 2020/37/B/ST1/01540. L.D. is affiliated with GNFM–INDAM (Istituto Nazionale di Alta Matematica) and acknowledges that this research is part of the EU Staff Exchange project 101086394 “Operator Algebras That One Can See”.